# Tactile SoftHand-A: 3D-Printed, Tactile, Highly-underactuated, Anthropomorphic Robot Hand with an Antagonistic Tendon Mechanism

### v1

@@ Abstract
For tendon-driven multi-fingered robotic hands, ensuring grasp adaptability while minimizing the number of actuators needed to provide human-like functionality is a challenging problem. Inspired by the Pisa/IIT SoftHand, this paper introduces a 3D-printed, highly-underactuated, five-finger robotic hand named the Tactile SoftHand-A, which features only two actuators. The dual-tendon design allows for the active control of specific (distal or proximal interphalangeal) joints to adjust the hand’s grasp gesture. We have also developed a new design of fully 3D-printed tactile sensor that requires no hand assembly and is printed directly as part of the robotic finger. This sensor is integrated into the fingertips and combined with the antagonistic tendon mechanism to develop a human-hand-guided tactile feedback grasping system. The system can actively mirror human hand gestures, adaptively stabilize grasp gestures upon contact, and adjust grasp gestures to prevent object movement after detecting slippage. Finally, we designed four different experiments to evaluate the novel fingers coupled with the antagonistic mechanism for controlling the robotic hand’s gestures, adaptive grasping ability, and human-hand-guided tactile feedback grasping capability. The experimental results demonstrate that the Tactile SoftHand-A can adaptively grasp objects of a wide range of shapes and automatically adjust its gripping gestures upon detecting contact and slippage. Overall, this study points the way towards a class of low-cost, accessible, 3D-printable, underactuated human-like robotic hands, and we openly release the designs to facilitate others to build upon this work. This work is Open-sourced at github.com/SoutheastWind/Tactile_SoftHand_A
@@ 1 Introduction
Anthropomorphic robotic hands can be used in various ways, such as for grasping/pick-and-place, remote teleoperation, autonomous in-hand manipulation and as prosthetics. However, developing economical and accessible anthropomorphic robotic hands that can approach the tactile sensibility and rich functionality of human hands has always been one of the most difficult and potentially rewarding challenges in robotics. The use of underactuated mechanisms is particularly appealing due to their simple structure, ease of maintenance, low cost, straightforward control, and adaptability to conform to the shape of objects (Piazza et al. (2019); Ozawa et al. (2014)). However, there are many such mechanisms for driving underactuated soft anthropomorphic hands, including pneumatic (Zhu et al. (2023), Lu et al. (2024)), tendon driven (Chen et al. (2018); Catalano et al. (2014); Li et al. (2022); Mizushima et al. (2018)) and linkage driven (Kim et al. (2020)).
@@
Figure 1: Exploded Front view (A) and back view (B) of the Tactile SoftHand-A. The main components shown include the driving tendons (in red), antagonistic tendons (in blue), differential mechanism due to the tendon layouts and couplings, driving and antagonistic motors, phalanges and tactile fingertips.
@@
Another important, but perhaps underappreciated, design aspect of anthropomorphic underactuated hands is the extension, or reset, of the finger joints, which plays a crucial role in the dexterity of the hand. Clearly, the human hand has different tendons to close and open the hand according to agonist and antagonist muscle groups. Fully actuated robotic hands realize the bending and reset of the joints through the forward and reverse rotation of the motor at the joint (Butterfaß et al. (2001)). However, almost all tendon-underactuated robot hands use passive elastic elements (Makino et al. (2017); Martin and Grossard (2014)) such as rubber bands (Catalano et al. (2014); Li et al. (2022); Ma and Dollar (2017)), torsion springs (Mnyusiwalla et al. (2016); Li et al. (2017); Jiang et al. (2014); Mizushima et al. (2018); Stuart et al. (2017)) or soft continuous materials (Mohd Faudzi et al. (2018)) to reset the finger joints (Ozawa et al. (2014)).
@@
Alongside the design and fabrication of the anthropomorphic hand, is its capability to be controlled, either autonomously or through human teleoperation. In this respect, having tactile sensing capabilities in the fingertips, and possibly over the rest of the hand, is necessary to provide direct information about contact. Historically, there have been many ways of integrating tactile fingertips into robot hands (Kappassov et al. (2015)); however, recently optical tactile sensors using cameras (Ward-Cherrier et al. (2018), Yuan et al. (2017)) have become prominent for various reasons, including the recent miniaturisation and cost-reduction in camera technology and the high-resolution nature of the sensory data fitting well with advances in computer vision. However, optical tactile sensors encounter bulky designs, making it difficult to integrate the componnent modules within the fingertip. Additionally, the complexity of the manufacturing process is another factor limiting the development of optical tactile sensors. Many such sensors, like DIGIT (Lambeta et al. (2020)) and 9DTact (Lin et al. (2023)), require specific molds and the casting of silicone to create soft skins. Others, like TacTip (Lepora (2021)), have involved injecting gel fillers into skin cavities. These manual manufacturing processes significantly restrict the shape design of optical tactile sensors, as well as the variety and distribution of inner structures such as markers. Furthermore, the manufacturing errors inherent in manual production can further impact the generalization performance of optical tactile sensors based on deep neural network methods.
This paper explores the design and development of an antagonist tendon mechanisms coupled with fully-integrated 3D-printed tactile sensors that together enhance the functionality of highly-underactuated anthropomorphic robotic hands. One significant advantage we find is that by coordinating the drive motor and antagonistic motor, it is possible to actively control the distal interphalangeal (DIP) or proximal interphalangeal (PIP) joints of the fingers, which can be coupled with the tactile sensing to thereby effectively control grasp gestures while maintaining the hand’s adaptive grasping. Another benefit is that active antagonism, by eliminating passive elastic elements, offers lower resistance during the hand-closing process, thereby increasing efficiency, responsiveness, and gripping strength.
Our main contributions in this work are:
1) We propose three different novel fingers with different antagonistic tendon mechanisms coupled with limited use of passive elastic components, which can achieve active control of the DIP and PIP joints respectively by using coordinating control of driving and antagonistic tendon.
2) We develop a fully 3D-printed tactile fingertip: utilizing advanced printing technology, we create multi-material and multi-hardness components in a single piece, eliminating the need for additional fabrication steps. Additionally, we developed a simple tactile model capable of detecting contact and slip. We are the first to introduce this fabrication because of a novel 3D-printing step we introduce that mixes clear support material with other materials to adjust their softness.
3) We build a novel, underactuated robot hand with antagonistic tendon, tactile fingertips and new differential mechanisms. We also developed a control scheme that integrates tactile sensing with the antagonistic tendon mechanism to optimize the functionality of this hand: the Tactile SoftHand-A, named because of its inspiration from the Pisa/IIT SoftHand (Catalano et al. (2014)). Its performance, including adaptability and grasping ability, is evaluated through finger and hand-synergy experiments, and tactile-controlled grasping experiments.
This paper is organized as follows: Section II briefly describes and compares several representative underactuated robotic hands, highlighting their main features and differences from the Tactile SoftHand-A; additionally, some works related to vision-based tactile sensors are introduced. Section III details the design of our novel fingers equipped with various antagonistic tendon mechanisms. Additionally, we introduce a physical model to elucidate the working principle and include relevant workspace simulations. Section IV describes the design and development of the TacTip fingertip, detailing the tactile model for detecting contact and slip. Section V discusses the development of the Tactile SoftHand-A based on the D-type finger design where the DIP joint is controllable. This section covers the overall design, differential mechanism, and tendon layout. Section VI details the control scheme of Tactile SoftHand-A. Sections VII and VIII introduce the experimental design and results, respectively. Discussion of the results and conclusions is presented in Sections IX and X.
Figure 2: The D-type finger where the antagonistic tendon terminates on the DIP joint. (A) Explored view. (B) Right view. (C) Left view. (D) Diagram showing the forces for analysis of the motion of the D-type fingertip.
2 Background and Related work
2.1 Tendon-driven Robot Hand
@@
Some near fully-actuated robotic hands seek to achieve complete dexterity, including the Shadow anthropomorphic hand, the UTAH-HIT hand (Jacobsen et al. (1986)) and the ROBISS hand (Mnyusiwalla et al. (2016)). The challenge with these hands is the complexity of the control needed to deploy the hand to grasp or manipulate objects and to operate the hand without damaging it.
@@
The Pisa/IIT SoftHand (Catalano et al. (2014)) is inspired by the form of the human hand, but seeks to simplify the control by using just one tendon through all joints that is driven by one actuator. This coupling between the elements of the hand gives it the ability to adapt to the shape of various objects. In addition, it uses a novel differential mechanism as an ingenious way to reduce the number of actuators while still having naturalistic hand movements that are characteristic of human hand motion (Sun et al. (2021); Xiong et al. (2016); Bicchi et al. (2011); Santello et al. (2016)). These differential mechanisms usually use the parallel slider mechanism (Sun et al. (2021)), moving-pulley mechanism (Gosselin et al. (2008); Gao et al. (2021); Chen et al. (2014)) or spring groups (Li et al. (2022)), which can be used to compose the synergy scheme (Catalano et al. (2014); Ozawa et al. (2014); Bicchi et al. (2011); Fan et al. (2018)).
@@
In the process of studying the above robot hands, we found that almost all of the underactuated tendon-driven hands use passive elements such as torsion springs (Mnyusiwalla et al. (2016); Li et al. (2017); Jiang et al. (2014); Mizushima et al. (2018)), high-strength springs (Makino et al. (2017); Martin and Grossard (2014); Stuart et al. (2017)) and elastic bands (Catalano et al. (2014); Li et al. (2022); Ma and Dollar (2017); Stuart et al. (2017)) to achieve joint reset motion. Few underactuated tendon-driven hands use active extension to reset the fingers. Of these, one has proposed a robotic finger with release tendon and grasp tendon (Crisman et al. (1996)). However, its bearing and tendon layout limits the rotation angle of its finger joint since the tendon can detach from the bearing or axis during rotation, so its motion space is relatively small (Crisman et al. (1996); Lee et al. (1994)). Here we use a different bearing layout to guide the flexion tendon and extension tendon to increase the joint rotation angle of the joint. Another study proposed a robotic finger based on the biomimetics of human finger muscles, with two actuators for the adductive motion of the joints and control of the extension of the finger through a third actuator (Shirafuji et al. (2014)). However, that finger has a complex tendon layout that is not conducive to the design and assembly of the overall hand. Furthermore, we have observed that currently, no tendon-driven robotic hand can decouple the DIP or PIP joint while ensuring adaptive capability, nor actively control the DIP or PIP joints to adjust the grasping gesture while maintaining adaptive grasping ability.
@@
Therefore, here we propose a novel tendon-driven finger design that achieves active extension and control of the DIP or PIP joints through coordinated control of driving and antagonistic tendons. Based on this new design, we have also developed a novel robotic hand that with just two motors can actively-manage closing, opening and grasping gestures.
2.2 Vision-based Tactile Sensing

### v2

@@ Abstract
A challenging and important problem for tendon-driven multi-fingered robotic hands is to ensure grasping adaptivity while minimizing the number of actuators needed to provide human-like functionality. Inspired by the Pisa/IIT SoftHand, this paper introduces a 3D-printed, highly-underactuated, tactile-sensorized, five-finger robotic hand named the Tactile SoftHand-A, which features an antagonistic mechanism to actively open and close the hand. Our proposed dual-tendon design gives options that allow active control of specific (distal or proximal interphalangeal) joints; for example, to adjust from an enclosing to fingertip grasp or to manipulate an object with a fingertip. We also develop and integrate a new design of fully 3D-printed vision-based tactile sensor within the fingers that requires minimal hand assembly. A control scheme based on analytically extracting contact location and slip from the tactile images is used to coordinate the antagonistic tendon mechanism (using a marker displacement density map, suitable for TacTip-based sensors). We perform extensive testing of a single finger, the entire hand, and the tactile capabilities to show the improvements in reactivity, load-bearing, and manipulability in comparison to a SoftHand that lacks the antagonistic mechanism. We also demonstrate the hand’s reactivity to contact disturbances including slip, and how this enables teleoperated control from human hand gestures. Overall, this study points the way towards a class of low-cost, accessible, 3D-printable, tactile, underactuated human-like robotic hands, and we openly release the designs to facilitate others to build upon this work. The designs are open-sourced at github.com/HaoranLi-Data/Tactile_SoftHand_A
@@ 1 Introduction
Anthropomorphic robotic hands can be used in various ways, such as for grasping/pick-and-place, remote teleoperation, autonomous in-hand manipulation, and as prosthetics. However, developing economical and accessible anthropomorphic robotic hands that can approach the tactile sensibility and rich functionality of human hands has always been one of the most difficult and potentially rewarding challenges in robotics. The use of underactuated mechanisms is particularly appealing due to their simple structure, ease of maintenance, low cost, straightforward control, and adaptability to conform to the shape of objects (Piazza et al. (2019); Ozawa et al. (2014)). To achieve this underactuation, a range of mechanisms have been proposed, including pneumatic (Zhu et al. (2023), Lu et al. (2025)), tendon-driven (Chen et al. (2018); Catalano et al. (2014); Li et al. (2022); Mizushima et al. (2018)), and linkage-driven (Kim et al. (2020)).
The article “A Century of Robotic Hands” (Piazza et al. (2019)) made some key observations about past trends and future directions in artificial hand design. One trend was towards a principled simplification of robotic hand design that maintains the advantages of an anthropomorphic structure while sensibly reducing the complexity of actuation and control. A second trend was an increased adoption of soft-robotics approaches, such as underactuated hands that utilize compliance or soft structures for adaptivity during grasping and manipulation. Both of these trends tend towards more biomimetic artificial hands, as human hands balance simplicity with functionality by making extensive use of compliance and underactuation.
@@
Figure 1: Exploded Front view (A) and back view (B) of the Tactile SoftHand-A. The main components shown include the agonist tendons (in red), antagonist tendons (in blue), differential mechanism due to the tendon layouts and couplings, agonist and antagonist motors, phalanges and tactile fingertips. Inside each joint is a connecting rod to maintain correct gear engagement. The overall length of Tactile SoftHand-A is 200 mm with palm width 90 mm and length 26 mm from palm to back cover.
The Pisa/IIT SoftHand (Catalano et al. (2014)) exemplifies these aspects of robot hand design: it is a highly-adaptive anthropomorphic tendon-driven hand that has naturalistic movements characteristic of human hand motion driven with just one motor. The hand is highly adaptive due to the soft robotic principle of adaptive synergies (Bicchi et al. (2011)) that synchronizes the motion of the fingers due to a common tendon looping through all the fingers (Catalano et al. (2014)). As a consequence, the hand exhibits dexterity comparable to far more highly-actuated hands, such as being able to grasp various shaped objects and even basic tool manipulations when the hand is mounted on a robot arm (Della Santina et al. (2017)). This dexterity is enhanced with tactile sensing in the fingertips (Ford et al. (2025)), enabling grasp-manipulation capabilities such as handling a flexible cup without crushing or dropping it under sudden changes in object weight, and being able to react to disturbances of held objects, e.g. in reaction to a human guiding the object motion.
An important direction to progress robotic hand research is to improve the dexterity of underactuated hands while retaining the benefits of simplified designs and eased control. In our view, the Pisa/IIT SoftHand gives a foundation for such progress, being anthropomorphic and exhibiting relatively high dexterity for a single motor. However, its original design is not easy to customize, for example to add actuators or to change the tendon routing; also, it is fabricated by injection molding. Inspired by the Yale GrabLab OpenHand series of 3D-printable customizable grippers (Ma and Dollar (2017); Odhner et al. (2014)), we designed the BPI (Bristol/Pisa/IIT) SoftHand that is entirely 3D-printed (Li et al. (2022)). For easier fabrication, the joints were simplified to pivot around linkage attachments, and soft synergies (springs) in the tendon-actuator coupling recovered some adaptivity and force transmission sacrificed by higher friction in the tendon routing due to the manufacturing.
@@
However, in researching highly-underactuated tendon-driven hands, we noticed that a key but underappreciated design aspect is the extension, or reset, of the finger joints. Clearly, the human hand has different tendons to close and open the hand according to agonist and antagonist muscle groups. Yet almost all tendon-underactuated robot hands use passive elastic elements (Makino et al. (2017); Martin and Grossard (2014); Salvietti et al. (2018)) such as rubber bands (Catalano et al. (2014); Li et al. (2022); Ma and Dollar (2017)), torsion springs (Mnyusiwalla et al. (2016); Li et al. (2017); Jiang et al. (2014); Mizushima et al. (2018); Stuart et al. (2017); Aukes et al. (2014); Chen et al. (2020)) or soft continuous materials (Mohd Faudzi et al. (2018)) to reset the finger joints (Ozawa et al. (2014)). As we will see later in this paper, an ever-present passive restoring force reduces the reactivity, dexterity, and load bearing of the hand.
This paper investigates the design and development of antagonistic tendon mechanisms within highly underactuated anthropomorphic robot hands. Furthermore, we take the view that the dexterity of the hand due to its actuation cannot be separated from that due to its control, which depends on the integration and deployment of tactile sensing. Therefore, we introduce an entirely 3D-printed, highly-underactuated, five-finger robotic hand named the Tactile SoftHand-A (Figure 1), which features two actuators: one each to open and close the hand. Within the hand structure, we introduce a new design of fully 3D-printed tactile sensor that requires minimal hand assembly and is printed directly within the robotic fingertip. We show that the combination of an antagonistic tendon mechanism and fingertip tactile sensing significantly improves the hand’s dexterity compared to the BPI SoftHand without these features.
An important improvement to the hand’s dexterity emerges from this antagonistic mechanism: it then becomes possible to actively control the distal interphalangeal (DIP) or proximal interphalangeal (PIP) joints of the fingers. This control is by coordinating the agonist and antagonist motors, as humans also seem to do when grasping delicate objects with our fingertips. When coupled with the tactile sensing, the antagonistic coordination can then effectively control grasp gestures while maintaining the hand’s adaptivity. There is a subtlety, however, that this additional dexterity depends upon the termination of the antagonist tendon and the limited use of passive elements. Another benefit is that active antagonism, by eliminating passive elastic elements, offers lower resistance during the hand-closing process, thereby increasing efficiency, responsiveness, and gripping strength.
Overall, our main contributions in this work are:
1) We introduce a class of antagonistic tendon mechanisms suitable for the fingers of highly-underactuated robot hands. These finger designs differ in the termination of the antagonist tendons, which determines whether the DIP or PIP joints can be actively controlled, respectively. Together, they constitute examples within a larger design space of finger structures and tendon mechanisms.
2) We develop a fully 3D-printed tactile fingertip that combines multi-material and multi-hardness components within the mechanical structures of the phalange, eliminating the need for additional fabrication steps. Additionally, we develop a simple tactile model capable of detecting contact and slip. We are the first to introduce this fabrication method because we introduce a novel 3D-printing step that mixes clear support with other materials to adjust their softness.
3) We combine the above to build a novel, underactuated robot hand with antagonistic tendons, tactile fingertips and new differential mechanisms: the Tactile SoftHand-A. We also develop a control scheme based on analytically extracting tactile features that integrates with the antagonistic tendon mechanism to deploy the functionality of this hand.
4) We evaluate the Tactile SoftHand-A’s performance, including adaptivity and grasping ability, through finger and hand-synergy experiments. We further evaluate the tactile performance, demonstrating the improved reactivity to contact disturbances, including slip, and how this enables synchronization with a human hand for use in teleoperation.
This paper is organized as follows. The next section briefly describes and compares several representative underactuated robotic hands, highlighting their main features and differences from the Tactile SoftHand-A; additionally, some works related to vision-based tactile sensors are introduced. The third section details the design of our novel fingers equipped with various antagonistic tendon mechanisms, additionally introducing a physical model of the working principle and a workspace simulation. The fourth section describes the design and development of the tactile fingertip and gives a simple analytic model for detecting contact and slip. The fifth section develops the Tactile SoftHand-A, covering the overall design with D-type finger, differential mechanism, and tendon layout. The sixth section details the control scheme, which is followed by two sections on the experimental design and results, covering the evaluation of a single finger, evaluation of the entire hand, grasping adaptivity, and tactile feedback control. Finally, we discuss these results and make our conclusions.
Figure 2: The D-type (distal) finger where the antagonist tendon terminates on the DIP joint. (A) Exploded view. (B) Right-side view. (C) Left-side view. (D) Diagram showing the forces for analysis of the motion of the D-type fingertip. Connecting rods within the joints in (A,D) are shown in yellow-brown. The finger’s total length, width, and height are 115 mm, 20 mm, and 23 mm, respectively.
2 Background and Related work
@@
2.1 Tendon-driven Robot Hands
@@
Some near fully-actuated robotic hands seek to achieve complete dexterity, including the Shadow anthropomorphic hand, M2 gripper (Ma et al. (2016)), the UTAH-HIT hand (Jacobsen et al. (1986)) and the ROBISS hand (Mnyusiwalla et al. (2016)). There are a variety of mechanisms that each have benefits and limitations. For example, the M2 Gripper utilizes a dual-tendon system to switch between underactuated and fully actuated modes, offering flexibility in grasping. However, it relies on passive reset mechanisms such as elastic bands or springs, which limit precise control of finger repositioning after releasing an object. Another approach is to adding electrostatic actuators to the joints of a tendon-driven robotic hand, to increase gripping force and dexterous ability by unlocking specific joints to decouple the finger joints (Aukes et al. (2014)). However, an agonist force is still needed to overcome the residual braking force and make the joint move. The challenge with all near fully-actuated hands is the complexity of the control needed to deploy the hand to grasp or manipulate objects without damaging it.
@@
The Pisa/IIT SoftHand (Catalano et al. (2014)) is inspired by the form of the human hand, but seeks to simplify the control by using just one tendon through all joints that is driven by one actuator. This coupling between the elements of the hand gives it the ability to adapt to the shape of various objects. In addition, it uses a novel differential mechanism as an ingenious way to reduce the number of actuators while still having naturalistic hand movements that are characteristic of human hand motion (Sun et al. (2021); Xiong et al. (2016); Bicchi et al. (2011); Santello et al. (2016)). These differential mechanisms usually use the parallel slider mechanism (Sun et al. (2021)), moving-pulley mechanism (Gosselin et al. (2008); Gao et al. (2021); Chen et al. (2014)), fixed pulley mechanism (Chen et al. (2020)) or spring groups (Li et al. (2022)), which can be used to compose the synergy scheme (Catalano et al. (2014); Ozawa et al. (2014); Bicchi et al. (2011); Fan et al. (2018)).
@@
In the process of studying the above robot hands, we found that almost all of the underactuated tendon-driven hands use passive elements such as torsion springs (Mnyusiwalla et al. (2016); Li et al. (2017); Jiang et al. (2014); Mizushima et al. (2018); Ma et al. (2016); Aukes et al. (2014)), high-strength springs (Makino et al. (2017); Martin and Grossard (2014); Stuart et al. (2017)) and elastic bands (Catalano et al. (2014); Li et al. (2022); Ma and Dollar (2017); Stuart et al. (2017)) to achieve joint reset motion. Few underactuated tendon-driven hands use active extension to reset the fingers. Of these, one has proposed a robotic finger with a release tendon and a grasp tendon (Crisman et al. (1996)). However, its bearing and tendon layout limits the rotation angle of its finger joint, since the tendon can detach from the bearing or axis during rotation, so its motion space is relatively small (Crisman et al. (1996); Lee et al. (1994)). Here, we use a different bearing layout to guide the flexion tendon and extension tendon to increase the joint rotation angle of the joint.
Another example of a robot hand employing a two-tendon mechanism for the robotic finger to enable active flexion and extension is by  Ruotolo et al. (2021). Additionally, the flexible connections between joints, coupled with the absence of mechanical limits, allow the fingers to bend in reverse, thereby increasing its operational workspace. However, a limitation remains: the joints are still coupled, and active control of the fingertip’s rotation cannot be achieved. Another study proposed a robotic finger based on the biomimetics of human finger muscles, with two actuators for the adductive motion of the joints and control of the extension of the finger through a third actuator (Shirafuji et al. (2014)). However, that finger has a complex tendon layout that is not conducive to the design and assembly of the overall hand. The fingers of the CLASH hand (Friedl and Roa (2020)) utilize two independent tendon systems to drive the MCP and DIP joints (without a PIP joint), thereby expanding the finger’s workspace. Additionally, a differential mechanism couples the DIP joints of two fingers, providing these joints with a degree of adaptability. However, while this design introduces adaptability between the DIP joints, it sacrifices the intrinsic adaptability between the MCP and DIP joints of the individual fingers.
It is also worth noting that in examples of more fully-actuated robot hands, both the DLR Hand II (Butterfaß et al. (2001)) and the Shadow Hand utilize modular fingers where the DIP joints are not independently controllable. To reduce the number of required motors, the DIP joints in these hands are coupled with the PIP joints; as a PIP joint flexes, the DIP joint follows proportionally. In contrast, the (fully-actuated) Allegro Hand achieves DIP joint control by incorporating an additional joint motor at the DIP joint. The BPI SoftHand-A, however, actively controls the DIP joint without extra motors, utilizing coordinated control between driving and antagonistic motors. This design enhances the hand’s dexterity and grasping capabilities.
Overall, we have observed that currently, no tendon-driven robotic hand can decouple the DIP or PIP joint while ensuring adaptive capability, nor actively control the DIP or PIP joints to adjust the grasping gesture while maintaining adaptive grasping ability.
@@
Therefore, here we propose a novel tendon-driven finger design that achieves active extension and control of the DIP or PIP joints through coordinated control of agonist and antagonist tendons. Based on this new design, we have also developed a novel robotic hand that with just two motors can actively manage closing, opening, and grasping gestures.
2.2 Tactile Sensing in Robot Hands
Alongside the design and fabrication of the anthropomorphic hand, is its capability to be controlled, either autonomously or through human teleoperation. In this respect, having tactile sensing capabilities in the fingertips, and possibly over the rest of the hand, is necessary to provide direct information about contact. Historically, there have been many ways of integrating tactile fingertips into robot hands (Kappassov et al. (2015)).
Compared to traditional electronic-based tactile sensors, vision-based tactile sensors (VBTS) using cameras have become prominent for various reasons, including the recent miniaturisation and cost reduction in camera technology and the high-resolution nature of the sensory data fitting well with advances in computer vision. The tactile images can provide rich physical information for robotic interaction tasks, such as texture features and object pose. Researchers have developed many VBTS, such as MIT’s GelSight (Yuan et al. (2017)), and BRL’s TacTip (Ward-Cherrier et al. (2018), Lepora (2021)) and BioTacTip (Li et al. (2024)).